why electron cant exist in nucleus with explanation

 

Grade:11

3 Answers

abhimanyu
5 years ago
as per uncertainty principle.. he electron has to exist in nucleus then it's speed should be more than speed than speed of light. but for that it has to be massless. but it do have mass.
Arun
5 years ago

I will explain you in easiest way.

(1) NUCLEAR SIZE : typically nuclei are less than 1 Fermi metre ( ≈10-¹⁴m approximately) in radius . If an electron exists inside a nucleus , the uncertainty in it's position (∆x ) may not exceed 10-¹⁴m . According to hinsenberg's uncertainty principle , uncertainty in electron moment is

∆p ≥ (h/2π) / ∆x

Putting the value of the constant we get

∆p ≥ 1.054 × 10 -³⁴ / 10-¹⁴

Solving we get ∆p ≥ 1.1× 10^-20 kg ms-¹

If this is the uncertainty in the electron's momentum , the momentum itself must be at least comparable in magnitude.

•: approximate momentum of the electron = p= 1.1 × 10 ^ -20 kg ms-¹. An electron whose momentum is 1.1×10^-20 kg ms-¹ has a K.E ( T) many times greater than it's rest energy m'c² .i.e »m’c² . Hence we can use the extreme relativistic formula.

T= (1.1×10^-20)× (3×10^8) = 3.3 × 10 -¹² J

Or 20.63 Mev

This shows that if an electron exists in the nucleus , the K.E. of the electron must be more than 20 Mev . Electron of such large energy are never found to be emitted during beta - decay . The maximum energy of a beta- particle emitted is only 2 to 3 Mev hence we can concluded that electron cannot be present within the nuclei.

Gayathri
5 years ago
Since according to einstein their are particles which have velocity  more than speed of light